# Strange stars admixed with dark matter: equiparticle model in a two fluid approach

### v2

@@ IV Strange stars admixed with dark matter
For the construction of strange stars through the EQP model admixed with fermionic or bosonic dark matter, all of them previously discussed, we consider the interfluid interaction uniquely coming from gravitational effects. For this purpose, the two-fluid formalism is a suitable treatment. In this approach, each fluid satisfies conservation of energy-momentum separately, that is equivalent to have $P(r)=P_{\mbox{\tiny vis}}(r)+P_{\mbox{\tiny DM}}(r)$ and $\mathcal{E}(r)=\mathcal{E}_{\mbox{\tiny vis}}(r)+\mathcal{E}_{\mbox{\tiny DM}}(r)$, with $r$ being the radial coordinate from the center of the star. This consideration leads to the following differential TOV equations to be solved,
@@
Technically, the procedure adopted to solve Eqs. (21)-(24) is the following. First we define the four initial conditions as $m_{\mbox{\tiny vis}}(0)=m_{\mbox{\tiny DM}}(0)=0$, $P_{\mbox{\tiny vis}}(0)=P^{c}_{\mbox{\tiny vis}}$, and $P_{\mbox{\tiny DM}}(0)=P^{c}_{\mbox{\tiny DM}}$, where $P^{c}_{\mbox{\tiny vis}}$ and $P^{c}_{\mbox{\tiny DM}}$ are the central pressures related to visible and dark matter, respectively, given by the equations of state presented in the previous sections. Then, for each set of initial conditions we use the fourth order Runge–Kutta method in order to obtain pressures and masses as functions of $r$. The radii $R_{\mbox{\tiny vis}}$ and $R_{\mbox{\tiny DM}}$ are defined as being the quantities that lead to $P_{\mbox{\tiny vis}}(R_{\mbox{\tiny vis}})/P^{c}_{\mbox{\tiny vis}}=0$ and $P_{\mbox{\tiny DM}}(R_{\mbox{\tiny DM}})/P^{c}_{\mbox{\tiny DM}}=0$, within a certain tolerance. These radii are also useful to determine $M_{\mbox{\tiny vis}}\equiv m_{\mbox{\tiny vis}}(R_{\mbox{\tiny vis}})$, and $M_{\mbox{\tiny DM}}\equiv m_{\mbox{\tiny DM}}(R_{\mbox{\tiny DM}})$. Therefore, the total mass of the respective star is $M=M_{\mbox{\tiny vis}}+M_{\mbox{\tiny DM}}$, and its radius is $R=R_{\mbox{\tiny vis}}$ if $R_{\mbox{\tiny vis}}>R_{\mbox{\tiny DM}}$, or $R=R_{\mbox{\tiny DM}}$ if $R_{\mbox{\tiny DM}}>R_{\mbox{\tiny vis}}$. This latter case identifies dark matter halo configurations for the star of mass $M$ and radius $R$. This mechanism is repeated for each point of the input equations of state, with the final collected pairs $(M,R)$ giving rise to the mass-radius diagram.
Here we impose that every star present a certain fraction of dark matter in its composition, i.e., a fixed value of $F_{\mbox{\tiny DM}}=M_{\mbox{\tiny DM}}/M$. In order to satisfy this restriction, we proceed as follows. First one needs to construct all the inputs for the TOV equations, namely, $P^{c}_{\mbox{\tiny vis}}$, $\mathcal{E}^{c}_{\mbox{\tiny vis}}$, $P^{c}_{\mbox{\tiny DM}}$ and $\mathcal{E}^{c}_{\mbox{\tiny DM}}$ (despite not directly used as initial conditions, the energy densities have also to be furnished since their relationships with the pressures are used to replace $\mathcal{E}$’s by $P$’s in the TOV equations). The visible matter equation of state (EOS) is straightforward, directly evaluated from Eqs. (11)-(12). For both DM models, fermionic and bosonic, we connect them with the EQP model by imposing that $\mathcal{E}_{\mbox{\tiny DM}}=f\mathcal{E}_{\mbox{\tiny vis}}$ where the energy density fraction, $f$, is a constant [71]. In the case of the bosonic DM model, this leads to an analytical expression for the dark boson density,
@@
For the fermionic dark matter case, on the other hand, there are not simple expressions relating the equations of states of quark and dark matter. The free parameters of the dark sector used here are $m_{\sigma}=15$ GeV [79], $C_{\sigma\phi}=0.1$ MeV-1 [79], $m_{\chi}=1.9$ GeV, and $C_{V}=3.26$ fm. The last two values, inside the ranges $0.5~{}\mbox{GeV}\leqslant m_{\chi}\leqslant 4.5~{}\mbox{GeV}$ [84] and $0.1~{}\mbox{fm}\leqslant C_{V}\leqslant 5~{}\mbox{fm}$ [71], were taken from [73]. As an illustration, we display BDM and FDM cases in Fig. 2 for fixed $f$.
@@
Figure 2: Dark matter pressure as a function of its respective energy density. Fermionic and bosonic cases.
Notice that for the chosen set of parameters the fermionic EOS is stiffer in comparison with the BDM one, as expected.
@@
Since the connection $\mathcal{E}_{\mbox{\tiny DM}}=f\mathcal{E}_{\mbox{\tiny vis}}$ is established, we solve the two-fluid TOV equations for huge set of values for $f$. For each solution related to a particular $f$, we evaluate $F_{\mbox{\tiny DM}}$ and select, among all the solutions, only the $(M,R)$ pairs that satisfy the particular value of $F_{\mbox{\tiny DM}}$ chosen. The final result consists in a mass-radius diagram representing quark stars, each one of them admixed with a fixed fraction of dark matter.
@@
As a remark, we emphasize that the two-fluid approach requires a detailed analysis on the stability of the final compact quark-DM star obtained, as the reader can verify in [85, 86]. In a one fluid system, the search for stable configurations consists in investigating the response of the star to radial oscillations in a Sturm-Liouville problem for the relative radial displacement and the pressure perturbation, both quantities depending on time through $e^{i\omega t}$ with $\omega$ being the eigenfrequency. Basically, the solutions of the new differential equations emerged from this investigation determines that stable stars are those in which $\omega^{2}>0$. Solutions presenting $\omega^{2}<0$ correspond to an exponential increasing of the perturbation, leading, consequently, to the collapse of the compact star. In the case of homogeneous stars, i.e., those presenting just one phase, the location of the last stable star ($\omega=0$) coincides with the maximum mass point in the mass-radius diagram, equivalent to the point in which $\partial M/\partial\mathcal{E}^{c}=0$.
@@
In the case of the two-fluid method, the generalization determines that at the onset of unstable configurations the number of visible ($N_{\mbox{\tiny vis}}$) and dark ($N_{\mbox{\tiny DM}}$) particles are stationary under variations of $\mathcal{E}^{c}_{\mbox{\tiny vis}}$ and $\mathcal{E}^{c}_{\mbox{\tiny DM}}$. As pointed out in [86, 85], it is analogous to diagonalize the matrix $\partial N_{i}/\partial\mathcal{E}^{c}_{j}$ ($i,j$=VIS, DM) with the aim of searching for its associated eingenvalues, claimed to be both positive for stable stars configurations. However, as discussed in [87], there is an indication that DM models presenting central pressures in the range of $P^{c}_{\mbox{\tiny DM}}\lesssim 5.1$ fm-4 generate stars in a stable region. Both DM models used in this work, bosonic and fermionic ones, are in agreement with this constraint. Furthermore, the authors of [88, 89] have used in their calculations, based on previous studies [90, 91], that the maxima of the fixed $F_{\mbox{\tiny DM}}$ curves represent the last stable stars in the mass-radius diagram, as it is in the case of one fluid systems. In our study we adopt the same procedure and exclude points beyond the maximum one. The final mass-radius diagrams for both models described here are shown in Fig. 3.
@@
Figure 3: Mass-radius relations for different fractions of dark matter in each strange stars. Results for (a) bosonic and (b) fermionic cases. The contours are related to data from the NICER mission, namely, PSR J0030+0451 [11, 12] and PSR J0740+6620 [13, 14], and the GW190425 event [10], all of them at $90\%$ credible level. The violet horizontal lines are also related to the PSR J0740+6620 pulsar [92].
@@
From this figure, one can verify that the maximum mass ($M_{\mbox{\tiny max}}$) of each diagram decreases as the mass fraction increases. This is a result compatible with previous studies that predict such a reduction in stars configurations in which a DM core is observed [78, 93, 94, 95, 96]. As a illustration, we plot $M_{\mbox{\tiny max}}$ versus $F_{\mbox{\tiny DM}}$ for both cases (BDM and FDM models) in Fig. 4.
@@
Figure 4: Maximum total mass versus mass fraction for the strange-DM stars presented in Fig. 3.
@@
It is worth noticing the strong linear relation between the maximum mass and the dark matter content. As a consequence, one has a constant decreasing rate of $M_{\mbox{\tiny max}}$ as a function of $F_{\mbox{\tiny DM}}$. Such behavior is hard to be predicted before calculations due to the nontrivial structure of both, the two-fluid equations and the equations of state used as input, in our case, the EQP model in the visible sector, and the fermionic/bosonic models in the description of dark matter. We also furnish the slope of the linear fitting obtained, namely, $-0.0343$ ($-0.315$) for bosonic (fermionic) model. Notice that both are very similar. Furthermore, these values are very close to $-0.0362$, result found in [97] in a study performed with a completely different quark model in the ordinary matter sector: the MIT bag model. For the dark sector, they have used a model made up of fermionic particles of mass $100$ GeV, number more than $50$ times greater than the value used in the current work for the FDM model. This result maybe indicates a possible universality in the (linear) decreasing behavior of the maximum mass in quark models admixed with dark matter.
@@
In Fig. 4 we can also verify that the FDM model produces maximum masses slightly greater than the ones generated from the BDM model, feature observed for the range of parameters investigated in our analysis. It suggests that DM equations of state are stiffer for the fermionic case with fixed dark matter content in comparison with the bosonic description of the dark sector as, indeed, we already shown in Fig. 2. This particular figure was plotted for a fixed $f$, but since the equations of state given as input are the same, the curves produced for a fixed value of $F_{\mbox{\tiny DM}}$ are also the same.
@@
Still regarding the features related to Fig. 3, we emphasize that the difference between the entire mass-radius curves generated from the quark-FDM approach and from the quark-BDM one are very small. We assign this effect to the dominance of the effective quark model over both, bosonic and fermionic DM models used in the analysis. In order to illustrate this statement, we display in Figs. 5 the profiles of the same star constructed by using both BDM and FDM models, with $M=1.81M_{\odot}$ and $10\%$ of dark matter in its composition.
@@
Figure 5: Pressure (panels a and b) and enclosed mass (panels c and d) as functions of radial coordinate for a mass fraction equal to $10\%$ and total stellar mass of $M=1.81M_{\odot}$. Full (dashed) lines: visible (dark) matter. For the bosonic (fermionic) case in panel a (b), the energy density fraction is $f=0.87$ ($0.6$).
@@
Notice from this example that the radius of the star is $R=R_{\mbox{\tiny vis}}\simeq 13.6$ ($13.7$) km, since $R_{\mbox{\tiny vis}}>R_{\mbox{\tiny DM}}$ in this case, by using the bosonic (fermionic) model in the dark sector. The main modification caused by the interchange of the DM model is the central pressure (at $r=0$) in both sectors. Nevertheless, $R_{\mbox{\tiny vis}}$ is still pratically equal for both models and determines the radius of the final star, according to the criterion used to define this quantity ($R$ corresponds to the outermost radius between $R_{\mbox{\tiny vis}}$ and $R_{\mbox{\tiny DM}}$). However, the situation is not the same for stars in which a DM halo is formed around it. We can verify such an effect for the case of $M=1.86M_{\odot}$ and $F_{\mbox{\tiny DM}}=10\%$, depicted in Fig. 6, for the strange star constructed with inclusion of bosonic dark matter in particular.
@@
Figure 6: Pressure as a function of radial coordinate for a mass fraction equal to $10\%$ and total stellar mass of $M=1.86M_{\odot}$. Full (dashed) lines: visible (dark) matter. Strange star admixed with bosonic dark matter with $m_{\sigma}=15$ GeV (panel a) and $m_{\sigma}=2$ GeV (panel b).
@@
In this figure we present two parametrizations of the BDM model, specifically with different values of the dark boson mass, namely, $m_{\sigma}=15$ GeV (standard value previously used) and $m_{\sigma}=2$ GeV. Notice that the decrease of the dark particle mass induces the situation of $R_{\mbox{\tiny DM}}>R_{\mbox{\tiny vis}}$ producing, consequently, the formation of a DM halo. In this case, the radius of the final star changes from $R=R_{\mbox{\tiny vis}}\simeq 13$ km to $R=R_{\mbox{\tiny DM}}\simeq 19$ km. Since fermionic and bosonic models generate similar mass-radius diagrams for $m_{\sigma}=15$ GeV and $m_{\chi}=1.9$ GeV, the decrease of $m_{\sigma}$ will produce a more pronounced difference between the diagrams obtained from FDM model with $m_{\chi}=1.9$ GeV in comparison with the ones determined from BDM model with $m_{\sigma}=2$ GeV, due to the emergence of DM halos in the latter system. These features are compatible with studies presented, for instance, in [96] where the authors used the two-fluid approach with bosonic DM, and in [98] where a fermionic DM model is applied in the description of stellar matter. In these works, the DM halo formation is observed when the dark particle (bosons or fermions) are decreased. In particular, since they treat neutron stars, hadronic models are used in the visible matter sector. In our work, on the other hand, strange stars are studied but the same effect is exhibited as well.
@@
Last but not least, we remark the compatibility of our results with the recent observational astrophysical data provided by the NICER mission regarding the millisecond pulsars PSR J0030+0451 [11, 12] and PSR J0740+6620 [13, 14], and with data from the gravitational wave event GW190425 [10] analyzed by LIGO and Virgo Collaboration. The same is observed with the PSR J0740+6620 data extracted from [92], corresponding to $M=(2.08\pm 0.07)M_{\odot}$ at 68.3% of credible level, and with the compact star named as HESS J1731-347 [16]. It points out to possible scenarios in which strange stars can admit dark matter in its interior.

### v3

@@ IV Strange stars admixed with dark matter
There are at least two main approaches used to describe strongly interacting matter admixed with dark matter at the mean-field level. One of them is based on the Higgs-portal models [91, 92], where DM is assumed to interact with ordinary particles through the exchange of the Higgs boson, with the whole system (DM + “visible sector”) described by a single Lagrangian density. It is shown that in hadronic models coupled to DM, this particular procedure leads to a simplification of the final equations of state, leading to a vanishing coupling between the distinct sectors (see [93] for instance). From the experimental point of view, such a result is corroborated by direct detection experiments [94, 95, 96, 97, 98], which only give upper bounds on the elastic scattering cross-section. Since DM particles and ordinary matter seem to be not interacting, it is reasonable to treat them as a two-fluid system in the interior of compact stars (quark, neutron, or hybrid stars, for instance); i.e., both sectors are coupled to each other only through gravity.
For the construction of strange stars through the EQP model admixed with fermionic or bosonic dark matter, all of them previously discussed, we consider the interfluid interaction uniquely coming from gravitational effects, as discussed above. For this purpose, the two-fluid formalism is used here. In this approach, each fluid satisfies the conservation of energy-momentum separately, which is equivalent to having $P(r)=P_{\mbox{\tiny vis}}(r)+P_{\mbox{\tiny DM}}(r)$ and $\mathcal{E}(r)=\mathcal{E}_{\mbox{\tiny vis}}(r)+\mathcal{E}_{\mbox{\tiny DM}}(r)$, with $r$ being the radial coordinate from the center of the star. This consideration leads to the following differential TOV equations to be solved,
@@
Technically, the procedure adopted to solve Eqs. (21)-(24) is the following: first, we define the four initial conditions as $m_{\mbox{\tiny vis}}(0)=m_{\mbox{\tiny DM}}(0)=0$, $P_{\mbox{\tiny vis}}(0)=P^{c}_{\mbox{\tiny vis}}$, and $P_{\mbox{\tiny DM}}(0)=P^{c}_{\mbox{\tiny DM}}$, where $P^{c}_{\mbox{\tiny vis}}$ and $P^{c}_{\mbox{\tiny DM}}$ are the central pressures related to visible and dark matter, respectively, given by the equations of state presented in the previous sections. Then, for each set of initial conditions we use the fourth-order Runge–Kutta method in order to obtain pressures and masses as functions of $r$. The radii $R_{\mbox{\tiny vis}}$ and $R_{\mbox{\tiny DM}}$ are defined as being the quantities that lead to $P_{\mbox{\tiny vis}}(R_{\mbox{\tiny vis}})/P^{c}_{\mbox{\tiny vis}}=0$ and $P_{\mbox{\tiny DM}}(R_{\mbox{\tiny DM}})/P^{c}_{\mbox{\tiny DM}}=0$, within a certain tolerance. These radii are also useful to determine $M_{\mbox{\tiny vis}}\equiv m_{\mbox{\tiny vis}}(R_{\mbox{\tiny vis}})$, and $M_{\mbox{\tiny DM}}\equiv m_{\mbox{\tiny DM}}(R_{\mbox{\tiny DM}})$. Therefore, the total mass of the respective star is $M=M_{\mbox{\tiny vis}}+M_{\mbox{\tiny DM}}$, and its radius is $R=R_{\mbox{\tiny vis}}$ if $R_{\mbox{\tiny vis}}>R_{\mbox{\tiny DM}}$, or $R=R_{\mbox{\tiny DM}}$ if $R_{\mbox{\tiny DM}}>R_{\mbox{\tiny vis}}$. This latter case identifies dark matter halo configurations for the star of mass $M$ and radius $R$. This mechanism is repeated for each point of the input equations of state, with the final collected pairs $(M,R)$ giving rise to the mass-radius diagram.
Here we impose that every star presents a certain fraction of dark matter in its composition, i.e., a fixed value of $F_{\mbox{\tiny DM}}=M_{\mbox{\tiny DM}}/M$. In order to satisfy this restriction, we proceed as follows. First, one needs to construct all the inputs for the TOV equations, namely $P^{c}_{\mbox{\tiny vis}}$, $\mathcal{E}^{c}_{\mbox{\tiny vis}}$, $P^{c}_{\mbox{\tiny DM}}$ and $\mathcal{E}^{c}_{\mbox{\tiny DM}}$ (although not directly used as initial conditions, the energy densities have also to be furnished since their relationships with the pressures are used to replace $\mathcal{E}$’s by $P$’s in the TOV equations). The visible matter equation of state (EOS) is straightforward, directly evaluated from Eqs. (11)-(12). For both fermionic and bosonic DM models, we connect them with the EQP model by imposing that $\mathcal{E}_{\mbox{\tiny DM}}=f\mathcal{E}_{\mbox{\tiny vis}}$ where the energy density fraction, $f$, is a constant [78]. In the case of the bosonic DM model, this leads to an analytical expression for the dark boson density,
@@
For the fermionic dark matter case, on the other hand, there are no simple expressions relating the equations of states of quark and dark matter. The free parameters of the dark sector used here are $m_{\sigma}=15$ GeV [86], $C_{\sigma\phi}=0.1$ MeV-1 [86], $m_{\chi}=1.9$ GeV, and $C_{V}=3.26$ fm. The last two values, inside the ranges $0.5~{}\mbox{GeV}\leqslant m_{\chi}\leqslant 4.5~{}\mbox{GeV}$ [80, 99] and $0.1~{}\mbox{fm}\leqslant C_{V}\leqslant 5~{}\mbox{fm}$ [80, 78], were taken from [80]. As an illustration, we display BDM and FDM cases in Fig. 3(a) for fixed $f$.
@@
Figure 3: Panel (a): dark matter pressure as a function of its respective energy density. Panel (b): dark matter sound speed. Fermionic and bosonic cases are represented by full and dashed lines, respectively.
Notice that for the chosen set of parameters, the fermionic EOS is stiffer in comparison with the BDM one, as expected. For the sake of completeness, we also display the sound speed related to the dark matter and defined as $v^{2}_{s\mbox{\tiny DM}}=\partial P_{\mbox{\tiny DM}}/\partial\mathcal{E}_{\mbox{\tiny DM}}$, see Fig. 3(b). From the figure, it is clear that the DM models considered here are causal, i.e., there is no superluminal equations of state in the dark sector.
@@
Since the connection $\mathcal{E}_{\mbox{\tiny DM}}=f\mathcal{E}_{\mbox{\tiny vis}}$ is established, we solve the two-fluid TOV equations for a huge set of values for $f$. For each solution related to a particular $f$, we evaluate $F_{\mbox{\tiny DM}}$ and select, among all the solutions, only the $(M,R)$ pairs that satisfy the particular value of $F_{\mbox{\tiny DM}}$ chosen. The final result consists of a mass-radius diagram representing quark stars, each one of them admixed with a fixed fraction of dark matter.
@@
As a remark, we emphasize that the two-fluid approach requires a detailed analysis of the stability of the final compact quark-DM star obtained, as the reader can verify in [100, 101]. In a one fluid system, the search for stable configurations consists of investigating the response of the star to radial oscillations in a Sturm-Liouville problem for the relative radial displacement and the pressure perturbation, both quantities depending on time through $e^{i\omega t}$ with $\omega$ being the eigenfrequency. Basically, the solutions of the new differential equations that emerged from this investigation determine that stable stars are those in which $\omega^{2}>0$. Solutions presenting $\omega^{2}<0$ correspond to an exponential increase of the perturbation, leading, consequently, to the collapse of the compact star. In the case of homogeneous stars, i.e., those presenting just one phase, the location of the last stable star ($\omega=0$) coincides with the maximum mass point in the mass-radius diagram, equivalent to the point in which $\partial M/\partial\mathcal{E}^{c}=0$.
In the case of the two-fluid method, the generalization determines that at the onset of unstable configurations the number of visible ($N_{\mbox{\tiny vis}}$) and dark ($N_{\mbox{\tiny DM}}$) particles are stationary under variations of $\mathcal{E}^{c}_{\mbox{\tiny vis}}$ and $\mathcal{E}^{c}_{\mbox{\tiny DM}}$, with $N_{\mbox{\tiny vis}}$ and $N_{\mbox{\tiny DM}}$ calculated from the differential equations given by
$\frac{dN_{\mbox{\tiny vis}}}{dr}$ $=\frac{4\pi r^{2}\rho_{\mbox{\tiny vis}}}{[1-2m(r)/r]^{1/2}},$ (27)
and
$\frac{dN_{\mbox{\tiny DM}}}{dr}$ $=\frac{4\pi r^{2}\rho_{\mbox{\tiny DM}}}{[1-2m(r)/r]^{1/2}},$ (28)
simultaneously coupled to the two-fluid TOV equations. As pointed out in [101, 100, 102, 103], the stability analysis is performed by investigating the diagonal version of the matrix $\partial N_{i}/\partial\mathcal{E}^{c}_{j}$ ($i,j$=VIS, DM), that comes from the equation
$\begin{pmatrix}\delta N_{\mbox{\tiny vis}}\\ \delta N_{\mbox{\tiny DM}}\end{pmatrix}=\begin{pmatrix}\partial N_{\mbox{\tiny vis}}/\partial\mathcal{E}^{c}_{\mbox{\tiny vis}}&\partial N_{\mbox{\tiny vis}}/\partial\mathcal{E}^{c}_{\mbox{\tiny DM}}\\ \partial N_{\mbox{\tiny DM}}/\partial\mathcal{E}^{c}_{\mbox{\tiny vis}}&\partial N_{\mbox{\tiny DM}}/\partial\mathcal{E}^{c}_{\mbox{\tiny DM}}\end{pmatrix}\begin{pmatrix}\delta\mathcal{E}^{c}_{\mbox{\tiny vis}}\\ \delta\mathcal{E}^{c}_{\mbox{\tiny DM}}\end{pmatrix}=0.$ (29)
The associate eigenvalues, named here as $\kappa_{1}$ and $\kappa_{2}$, are claimed to satisfy the condition
$\kappa_{1}>0,\qquad\kappa_{2}>0$ (30)
@@
for stable stars’ configurations. However, as discussed in [104], there is an indication that DM models presenting central pressures in the range of $P^{c}_{\mbox{\tiny DM}}\lesssim 5.1$ fm-4 generate stars in a stable region. Both DM models used in this work, bosonic and fermionic ones, are in agreement with this constraint. Furthermore, the authors of [105, 106] have used in their calculations, based on previous studies [107, 108], that the maxima of the fixed $F_{\mbox{\tiny DM}}$ curves represent the last stable stars in the mass-radius diagram, as it is in the case of one fluid systems. Here we present in Fig. 4 the final mass-radius diagrams for both quark-DM models previously described. We also mark with circles the last stable stars ($F_{\mbox{\tiny DM}}>0$) obtained from the implementation of the condition given in Eq. (30).
@@
Figure 4: Mass-radius relations for different fractions of dark matter in each strange stars. Results for (a) bosonic and (b) fermionic cases. It also plotted data related to the NICER mission concerning the pulsars PSR J0030+0451 (dashed [15] and solid [16] brown contours), and PSR J0740+6620 (solid indigo countour [20] and indigo rectangular band [21]). We also compared our results with data from the GW190425 event [14] (turquoise contours), and the compact object named HESS J1731-347 [23] (orange contours). The circles represent the last stable stars obtained from the condition given in Eq. (30).
It is worth noticing that, as expected, for small values of $F_{\mbox{\tiny DM}}$, the last stable star ($M_{ls}$) is not so different from the star with maximum mass ($M_{\max}$). The extreme case is the curve related to $F_{\mbox{\tiny DM}}=0$ (no DM included) for which $M_{ls}=M_{\max}$.
@@
As another feature presented in the figure, one can verify that $M_{ls}$ of each diagram decreases as the mass fraction increases. This is a result compatible with previous studies that predict such a reduction in star configurations in which a DM core is observed [85, 109, 110, 111, 112]. As an illustration, we plot $M_{ls}$ versus $F_{\mbox{\tiny DM}}$ for both cases (BDM and FDM models) in Fig. 5.
@@
Figure 5: Mass of the last stable star, $M_{ls}$, versus mass fraction for the strange-DM stars represented by circles in Fig. 4.
@@
It is worth noticing the strong linear relation between the mass of the last stable star and the dark matter content. As a consequence, one has a constant decreasing rate of $M_{ls}$ as a function of $F_{\mbox{\tiny DM}}$. Such behavior is hard to predict before calculations due to the nontrivial structure of both, the two-fluid equations and the equations of state used as input, in our case, the EQP model in the visible sector, and the fermionic/bosonic models in the description of dark matter. We also furnish the slope of the linear fitting obtained, namely, $-0.04014$ ($-0.03489$) for the bosonic (fermionic) model. Notice that both are very similar. Furthermore, these values are very close to $-0.0362$, the result found in [113] in a study performed with a completely different quark model in the ordinary matter sector: the MIT bag model. For the dark sector, they have used a model made up of fermionic particles of mass $100$ GeV, a number more than $50$ times greater than the value used in the current work for the FDM model. This result may indicate a possible universality in the (linear) decreasing behavior of the maximum mass in quark models admixed with dark matter.
@@
In Fig. 5 we can also verify that the FDM model produces $M_{ls}$ slightly greater than the ones generated from the BDM model, a feature observed for the range of parameters investigated in our analysis. It suggests that DM equations of state are stiffer for the fermionic case with fixed dark matter content in comparison with the bosonic description of the dark sector as, indeed, we already shown in Fig. 3. This particular figure was plotted for a fixed $f$, but since the equations of state given as input are the same, the curves produced for a fixed value of $F_{\mbox{\tiny DM}}$ are also the same.
@@
Still, regarding the features related to Fig. 4, we emphasize that the difference between the entire mass-radius curves generated from the quark-FDM approach and from the quark-BDM one is very small. We assign this effect to the dominance of the effective quark model over both, bosonic and fermionic DM models used in the analysis. In order to illustrate this statement, we display in Figs. 6 the profiles of the same star constructed by using both BDM and FDM models, with $M=1.81M_{\odot}$ and $10\%$ of dark matter in its composition.
@@
Figure 6: Pressure (panels a and b) and enclosed mass (panels c and d) as functions of the radial coordinate for a mass fraction equal to $10\%$ and total stellar mass of $M=1.81M_{\odot}$. Full (dashed) lines: visible (dark) matter. For the bosonic (fermionic) case in panel a (b), the energy density fraction is $f=0.87$ ($0.6$).
@@
Notice from this example that the radius of the star is $R=R_{\mbox{\tiny vis}}\simeq 13.6$ ($13.7$) km, since $R_{\mbox{\tiny vis}}>R_{\mbox{\tiny DM}}$ in this case, by using the bosonic (fermionic) model in the dark sector. The main modification caused by the interchange of the DM model is the central pressure (at $r=0$) in both sectors. Nevertheless, $R_{\mbox{\tiny vis}}$ is still practically equal for both models and determines the radius of the final star, according to the criterion used to define this quantity ($R$ corresponds to the outermost radius between $R_{\mbox{\tiny vis}}$ and $R_{\mbox{\tiny DM}}$). However, the situation is not the same for stars in which a DM halo is formed around it. We can verify such an effect for the case of $M=1.59M_{\odot}$ and $F_{\mbox{\tiny DM}}=10\%$, depicted in Fig. 7, for the strange star constructed with the inclusion of bosonic dark matter in particular.
@@
Figure 7: Pressure as a function of the radial coordinate for a mass fraction equal to $10\%$ and total stellar mass of $M=1.59M_{\odot}$. Full (dashed) lines: visible (dark) matter. Strange star admixed with the bosonic dark matter with $m_{\sigma}=15$ GeV (panel a) and $m_{\sigma}=2$ GeV (panel b).
@@
In this figure, we present two parametrizations of the BDM model, specifically with different values of the dark boson mass, namely, $m_{\sigma}=15$ GeV (standard value previously used) and $m_{\sigma}=2$ GeV. Notice that the decrease of the dark particle mass induces the situation of $R_{\mbox{\tiny DM}}>R_{\mbox{\tiny vis}}$ producing, consequently, the formation of a DM halo. In this case, the radius of the final star changes from $R=R_{\mbox{\tiny vis}}\simeq 14$ km to $R=R_{\mbox{\tiny DM}}\simeq 19$ km. Since fermionic and bosonic models generate similar mass-radius diagrams for $m_{\sigma}=15$ GeV and $m_{\chi}=1.9$ GeV, the decrease of $m_{\sigma}$ will produce a more pronounced difference between the diagrams obtained from FDM model with $m_{\chi}=1.9$ GeV in comparison with the ones determined from BDM model with $m_{\sigma}=2$ GeV, due to the emergence of DM halos in the latter system. These features are compatible with studies presented, for instance, in [112] where the authors used the two-fluid approach with bosonic DM and in [114] where a fermionic DM model is applied in the description of stellar matter. In these works, the DM halo formation is observed when the dark particles (bosons or fermions) are decreased. In particular, since they treat neutron stars, hadronic models are used in the visible matter sector. In our work, on the other hand, strange stars are studied but the same effect is exhibited as well.
We also emphasize some astrophysical consequences of the findings presented in our work. Recently, some studies pointed out the impact of DM on the cooling process of neutron stars, as one can see in [115, 116] for instance. The authors of [115] conclude that, despite being thermally inert, DM indirectly impacts the thermal evolution of neutron stars. In that case, the one-fluid approach is adopted, and the study shows that admixed DM systems with higher values of dark Fermi momentum (dark matter content) cause an acceleration of thermal evolution, namely, a more rapid cooling of the star. In Ref. [116], on the other hand, the two-fluid formalism is used, but the same qualitative feature is found, i.e., an increase in the DM fraction $F_{\mbox{\tiny DM}}$ induces more efficient and rapid cooling due to the earlier onset of the Direct Urca process. The effects of DM on the cooling of strange stars (two-fluid formalism) could also be studied in future work to verify whether the previously observed features also arise in these compact objects.
In particular, it could help determine whether the HESS J1731-347 measurement [23] can be used as a possible signature of strange stars, as proposed in [117, 118, 119, 120, 121, 122, 123, 124]. Notice that the strange stars presented here are compatible with these data, regardless of the DM content. However, unpaired quark matter could lead to difficulties in reproducing the surface temperature of this particular compact object, according to [121].
@@
Last but not least, we remark the compatibility of our results with the recent observational astrophysical data provided by the NICER mission regarding the millisecond pulsars PSR J0030+0451 [16, 15] and PSR J0740+6620, and with data from the gravitational wave event GW190425 [14] analyzed by LIGO and Virgo Collaboration. With regard to the PSR J0030+0451 pulsar, of mass, corresponding to $M=(2.08\pm 0.07)M_{\odot}$ [17], we have included updated data analyzed by two different groups. The first one used data from 3.6 years of observation (from 2018 September to 2022 April) [20], and the second one, performed in Ref. [21], incorporated previous NICER and XMM-Newton observations along with additional NICER data of about $1.1$ Ms of observation in comparison with previous studies [19, 18]. In general, our results point out possible scenarios in which strange stars can admit dark matter in their interior.